### v2

# On the Limits of Self-Improving in Large Language Models: The Singularity Is Not Near Without Symbolic Model Synthesis

Abstract
We formalise recursive self-training in Large Language Models (LLMs) and Generative AI as a discrete-time dynamical system. We prove that if the proportion of exogenous, externally grounded signal $\alpha_{t}$ vanishes asymptotically ($\alpha_{t}\to 0$), the system undergoes degenerative dynamics. We derive two fundamental failure modes: (1) Entropy Decay, where finite sampling effects induce monotonic loss of distributional diversity, and (2) Variance Amplification, where the absence of persistent grounding causes distributional drift via a random-walk mechanism. These behaviours are architectural invariants of distributional learning on finite samples. We show that the collapse results apply specifically to closed-loop density matching without persistent external signal. Systems with non-vanishing exogenous grounding fall outside this regime. However, mainstream Singularity, AGI, and ASI narratives typically posit systems that become increasingly autonomous and require little to no human or external intervention for self-improvement. In that autonomy regime, the vanishing-signal condition is satisfied, and collapse follows under KL-based objectives. To overcome these limits, we propose neurosymbolic integration based on algorithmic probability and program synthesis. The Coding Theorem Method (CTM) enables identification of generative mechanisms rather than mere correlations, escaping the distribution-only constraints that bind standard statistical learning. We conclude that fully autonomous recursive density matching leads to degenerative fixed points, whereas externally anchored or mechanism-based approaches operate under fundamentally different asymptotic dynamics.
Keywords: Large Language Models (LLMs), Model Collapse, Recursive Self-Improvement, Entropy Decay, Coding Theorem Method, Algorithmic Information Dynamics.
1 Introduction
The notion of a technological or AI Singularity, popularised by Vernor Vinge and Ray Kurzweil (14), posits a future inflection point where artificial intelligence surpasses human intellect, leading to an “intelligence explosion” of unforeseeable consequence (25; 15). Central to this hypothesis is the concept of recursive self-improvement: an AI system with the capacity to inspect and enhance its own architecture or training processes would initiate a positive feedback loop, with each generation of the AI being more intelligent than the last, leading to exponential growth in its capabilities.
The recent and remarkable successes of Generative Artificial Intelligence (GenAI), particularly Large Language Models (LLMs) like GPT-5 (17) and Diffusion Models for image synthesis (19), have reignited speculation about the proximity of this event. These models demonstrate an unprecedented ability to generate fluent text, create photorealistic images, and synthesise complex data, leading some to believe they are foundational steps towards Artificial General Intelligence (AGI). The assumption is that by scaling these models and enabling them to learn from the vast quantities of data they can generate, we might trigger the prophesied recursive improvement cycle.
This paper challenges this assumption directly. We argue that the very mechanism proposed for self-improvement–training on self-generated data–is, in fact, a pathway to self-destruction. This phenomenon, empirically observed and termed model collapse or the ‘curse of recursion’ (21), describes the progressive degradation of a model’s performance as its training data becomes increasingly polluted with its own synthetic outputs. Rather than ascending towards superintelligence, the model’s internal representation of the world contracts and distorts, converging towards a degenerate state of low diversity and high bias.
Our contribution is to move beyond empirical observation and provide a formal mathematical proof of the inevitability of model collapse. We model the self-referential training process as a dynamical system on the space of probability distributions and demonstrate that, under the condition of a diminishing supply of fresh, authentic data, this system is guaranteed to converge to a fixed point that is a distorted and impoverished version of the true data distribution. This conclusion is robust and applies not only to single LLMs but also to complex ecosystems of interacting models and multi-modal systems.
Furthermore, we contend that this mathematical limitation reflects a deeper philosophical boundary. Drawing upon Immanuel Kant’s distinction between analytic and synthetic judgements (12), we argue that current GenAI is fundamentally an analytic engine. It excels at analysing, recombining, and interpolating the vast patterns contained within its human-generated training data. It cannot, however, generate synthetic knowledge–truly novel concepts, laws, or truths that are not simply derivative of its input. The Singularity requires a capacity for synthetic knowledge generation, which is absent in the current paradigm.
Recent empirical work on model collapse has clarified that degenerative dynamics depend critically on the proportion of synthetic to real data. While early discussions considered scenarios in which synthetic data progressively replace real data, more recent studies show that collapse can be mitigated when real data continue to accumulate at a sufficient rate (2). In the formalism developed in this paper, this corresponds to maintaining $\inf_{t}\alpha_{t}>0$, where $\alpha_{t}$ denotes the proportion of exogenous, externally grounded signal.
The central question for Singularity, AGI, and ASI arguments is therefore not whether collapse can be avoided under persistent external grounding, but whether recursive self-improvement is assumed to become increasingly autonomous. Strong versions of these hypotheses typically posit systems that require progressively less human or environmental input in order to improve themselves. Such autonomy implies that the relative contribution of externally grounded signal becomes negligible over time, i.e. that $\alpha_{t}\to 0$. It is precisely this autonomy regime that our theorems analyse.
2 Preliminaries and Notation
This section fixes all notation and technical terms used in the manuscript. We separate (i) distributional objects and learning operators, (ii) information-theoretic quantities and inequalities, and (iii) algorithmic information estimators and perturbation-based mechanism analysis.
For a discrete distribution $P$ on $\mathcal{X}$, Shannon entropy is $H(P)=-\sum_{x\in\mathcal{X}}P(x)\log P(x)$. We denote the Kullback–Leibler divergence by $D_{KL}(P\|Q)=\sum_{x}P(x)\log\frac{P(x)}{Q(x)}$.
In realistic training scenarios, the model does not have access to the full distribution $Q_{t}$ or $P$, but only to a finite dataset $\mathcal{D}_{t}=\{x_{1},\dots,x_{N}\}$ of size $N$ sampled i.i.d. from the source distribution. We denote the empirical distribution formed by these samples as $\widehat{Q}_{t}$.
To analyse dynamics in high-dimensional spaces (like the latent space of an LLM), we assume there exists a feature map $\phi:\mathcal{X}\to\mathbb{R}^{d}$ (e.g., the transformer embedding) and define the first moment (mean) of the distribution as $\mu_{Q}=\mathbb{E}_{x\sim Q}[\phi(x)]$. We denote the model update as an operator $\mathcal{T}$ that maps a target distribution to a model parameterisation, minimising divergence:
$Q_{t+1}\;=\;\arg\min_{Q\in\mathcal{Q}}D_{KL}(\widehat{P}^{\prime}_{t}\,\|\,Q),$ (1)
where $\widehat{P}^{\prime}_{t}$ is the empirical mixture derived from finite samples.
Throughout, $I(\cdot;\cdot)$ denotes Shannon mutual information. For discrete random variables $U,V$ it is
$I(U;V)\;=\;\sum_{u,v}p(u,v)\log\frac{p(u,v)}{p(u)p(v)}.$ (2)
We use mutual information only to formalise limitations of distribution-only learning pipelines. For a Markov chain $M\to X\to Y$, the data-processing inequality (DPI) states
$I(M;Y)\;\leq\;I(M;X).$ (3)
We work with iterative operators that admit one-step bounds of the form
$D_{f}^{(t+1)}\;\leq\;c\,D_{f}^{(t)}+\delta,\qquad c\in(0,1],\ \delta\geq 0,$ (4)
where $c$ is a contraction factor and $\delta$ captures approximation error. Iterating yields
$D_{f}^{(n)}\;\leq\;c^{n}D_{f}^{(0)}\;+\;\sum_{i=0}^{n-1}c^{i}\delta,$ (5)
so smaller $c$ corresponds to stronger correction per iteration, up to an error floor induced by $\delta$.
2.1 Coding and Block Decomposition Methods and Algorithmic Information Dynamics
Kolmogorov complexity and algorithmic probability provide a formal framework for measuring information in terms of generative mechanisms rather than statistical frequencies. Given a fixed universal Turing machine, the Kolmogorov complexity of a finite object is defined as the length of the shortest programme that produces that object as output. It therefore quantifies information as minimal description length in programme space. An object is considered algorithmically simple if it admits a short generating programme, and algorithmically random if no description significantly shorter than the object itself exists.
Algorithmic probability complements this notion by assigning to each object the probability that a randomly generated programme will produce it. Short programmes contribute more weight than long ones, yielding a universal prior that favours simpler generative explanations. The coding theorem links the two concepts: objects with high algorithmic probability have low Kolmogorov complexity, and vice versa. Together, these notions shift the focus from observed statistical regularities to the underlying mechanisms capable of generating the data, forming the basis for a mechanistic interpretation of information beyond Shannon entropy.
Fix a universal prefix Turing machine $U$. For a computable object $o$, its Kolmogorov complexity is $K(o)=\min\{|p|:U(p)=o\}$. The algorithmic probability is $m(o)=\sum_{p:U(p)=o}2^{-|p|}$. The Coding Theorem connects them: $\log m(o)=K(o)+O(1)$.
The Coding Theorem Method (CTM) approximates algorithmic probability by enumerating a reference class $\mathcal{M}$ of small Turing machines:
$\widehat{m}_{\mathrm{CTM}}(o)\;=\;\frac{1}{|\mathcal{M}|}\sum_{M\in\mathcal{M}}\mathbf{1}\{U_{M}\downarrow=o\},\quad\mathrm{CTM}(o)=-\log\widehat{m}_{\mathrm{CTM}}(o).$ (6)
To scale beyond the small-object regime, BDM decomposes an object $o$ into blocks of size $k$:
$\mathrm{BDM}_{k}(o)\;=\;\sum_{i}(\mathrm{CTM}(b_{i})+\log n_{i}),$ (7)
where $n_{i}$ is the multiplicity of block $b_{i}$.
Algorithmic Information Dynamics (AID) is a theoretical framework that studies the structure, evolution and causal content of systems through the lens of Algorithmic Information Theory. Rather than relying on statistical regularities, it characterises patterns, randomness and complexity in terms of the length of the shortest effective description capable of generating an object or process.
The central idea is that the behaviour of a system can be understood by analysing how its algorithmic information content changes under perturbations. By estimating contributions of individual elements to the overall description length, the framework aims to distinguish structural, causal components from incidental or noisy ones. In this way, Algorithmic Information Dynamics provides tools for inferring generative mechanisms, identifying causal pathways and quantifying the directionality of information flow in deterministic and stochastic systems alike.
Although rooted in uncomputable quantities such as Kolmogorov complexity, the framework employs computable approximations and decomposition methods to make empirical analysis possible across domains including dynamical systems, networks and biological processes.
AID quantifies the algorithmic causal effect of a perturbation $\tau$ by the change in complexity:
$\Delta_{\tau}(o)\;=\;\mathrm{BDM}_{k}(\tau(o))-\mathrm{BDM}_{k}(o).$ (8)
2.2 Neurosymbolic operators
A neurosymbolic operator is a transformation that combines two distinct modes of inference:
  • Subsymbolic / statistical learning, typically implemented by neural networks that approximate probability distributions from data.
  • Symbolic / structural reasoning, which operates over discrete representations such as rules, equations, graphs, logical constraints, or programmes.
A neurosymbolic operator refers to a transformation that integrates statistical learning with structured, rule-based or mechanistic reasoning within a single update process. Modern neural systems are highly effective at modelling complex probability distributions from large datasets, capturing high-dimensional correlations and producing flexible interpolations. However, such systems operate primarily in distribution space: they adjust parameters to minimise a loss function defined over observed frequencies, without explicit access to underlying structural constraints or generative mechanisms.
Symbolic systems, by contrast, manipulate discrete representations such as logical rules, equations, causal graphs, or programmes. They operate in structure space, enforcing invariants, compositional rules, and mechanistic relations that are not reducible to correlation alone. While neural models excel at approximation, symbolic systems excel at constraint, explanation, and generalisation under transformation.
A neurosymbolic operator combines these two regimes. Formally, it updates a learned model not only to improve statistical fit to data, but also to satisfy or approximate a set of structural constraints derived from symbolic representations. The operator therefore acts simultaneously in distribution space and in structure space. In practice, this may take the form of projecting a neural model onto a constraint set defined by logical rules, causal relations, conservation laws, or syntactic grammars, or of guiding model updates through programme or equation synthesis consistent with the data.
In an algorithmic framework, the notion extends further: the symbolic component may consist of candidate generative programmes weighted by algorithmic probability, and updates are evaluated not solely by reduction in statistical divergence but by their coherence under structured perturbations. In this setting, a neurosymbolic operator is not merely a hybrid architecture but a principled learning mechanism that integrates statistical evidence with mechanistic structure, thereby moving beyond purely correlational inference.
The intuition is that neural models are powerful at interpolation and pattern extraction in high-dimensional spaces, while symbolic systems are powerful at enforcing invariants, causal structure, compositionality, and mechanistic constraints. A neurosymbolic operator acts as a bridge: it updates or transforms a model by jointly optimising statistical fit and structural consistency.
We can define a one-step update as a composition of operators:
$Q_{t+1}\;=\;\mathcal{T}_{\alpha_{t}}\circ\mathcal{C}_{t}\circ\Pi_{\mathcal{S}}(Q_{t}),$
where $\Pi_{\mathcal{S}}$ is symbolic projection, $\mathcal{C}_{t}$ is causal correction, and $\mathcal{T}_{\alpha_{t}}$ is statistical fitting.
3 Background and Related Work
The modern conception of the technological Singularity is rooted in the idea of an “intelligence explosion,” first detailed by I. J. Good (9). Good argued that an “ultraintelligent machine”–defined as a machine that can far surpass all the intellectual activities of any man however clever–would be the last invention that man need ever make, provided that the machine is docile enough to tell us how to keep it under control. The core of the logic is recursive: such a machine could design even better machines, which would in turn design even better ones, creating a positive feedback loop of rapidly accelerating intelligence.
This idea requires two key preconditions: (1) the existence of an AI that is intelligent enough to understand and modify its own source code or training methodology and (2) that these modifications lead to a consistent increase in its own general intelligence. Proponents of the Singularity, such as (15), view the exponential growth in computing power (Moore’s Law) as a direct trajectory towards this point. However, this view often conflates computational capacity with the architectural leaps required for genuine intelligence growth. The core question is not whether a machine can be made more powerful, but whether it can make itself more intelligent in a meaningful and unbounded way.
3.1 Generative AI as Distribution Learners
Modern GenAI models, irrespective of their modality, share a common mathematical foundation: they are designed to learn and sample from a complex, high-dimensional probability distribution.
Let $\mathcal{X}$ be a data space (e.g., the space of all possible images, texts, or protein structures). We assume there exists a true, underlying data distribution $P(x)$ for $x\in\mathcal{X}$, from which real-world data is sampled. A generative model, parametrised by $\theta\in\Theta$, aims to learn an approximation of this distribution, denoted $Q_{\theta}(x)$.
  • Autoregressive Models (e.g., LLMs): For a sequence of tokens $x=(x_{1},\dots,x_{L})$, these models learn the conditional probability of the next token given the preceding ones. The joint probability is factorised as $Q_{\theta}(x)=\prod_{i=1}^{L}Q_{\theta}(x_{i}\mid x_{<i})$. Training typically involves minimising the negative log-likelihood (cross-entropy) on a large corpus of text, which is equivalent to minimising the Kullback-Leibler (KL) divergence $D_{KL}(P||Q_{\theta})$.
  • Diffusion Models: These models learn to reverse a diffusion process that gradually adds noise to data. They define a sequence of latent variables that corrupt an initial data point $x_{0}\sim P(x)$ into pure Gaussian noise $x_{T}$. The model then learns the reverse process $Q_{\theta}(x_{t-1}\mid x_{t})$, allowing it to generate a sample $x_{0}$ by starting with random noise and iteratively denoising it. This process implicitly defines a complex distribution $Q_{\theta}(x)$.
  • Generative Adversarial Networks (GANs): GANs (10) use a two-player game between a generator $G$ that produces samples and a discriminator $D$ that tries to distinguish them from real samples. The generator $G$ implicitly defines the model distribution $Q_{\theta}(x)$. The feedback from the discriminator guides the generator to produce samples that are increasingly indistinguishable from the true data distribution $P(x)$.
In all cases, the goal is to make $Q_{\theta}(x)$ as close as possible to $P(x)$. The Singularity hypothesis, when applied to these models, implies that a model $M_{t}$ with distribution $Q_{t}$ could generate data to train a successor model $M_{t+1}$ with distribution $Q_{t+1}$ such that $Q_{t+1}$ is a “better” approximation of $P$ or represents a “more intelligent” distribution. Our paper will show this is not the case.
3.2 Model Collapse: The Curse of Degenerative Recursion
The theoretical foundation for our argument rests on the growing body of literature concerning model collapse. (21) provided a seminal empirical and theoretical study, showing that learning from data generated by other models causes a feedback loop that makes models “forget” the true underlying distribution. They demonstrate this for Gaussian Mixture Models and show empirically for LLMs and diffusion models that diversity is rapidly lost, with the tails of the original distribution being forgotten first.
This phenomenon is not new, though its implications for the Singularity are underexplored. In GAN literature, a similar issue known as “mode collapse” occurs when the generator learns to produce only a few distinct types of samples that can fool the discriminator, failing to capture the full diversity of the data distribution (4). The work in (3) refers to this as “self-consuming” loops, warning that the proliferation of synthetic data on the internet could “contaminate” the training data for future models, leading to a gradual decay in their quality.
Previous studies on model collapse demonstrate empirically that generative models degrade when trained on synthetic data. By contrast, our contribution is not to show that collapse occurs, but to prove that it must occur under the autonomy condition required for recursive self-improvement. We formalise self-training as a dynamical system and show that, as external grounding vanishes, the system converges to a degenerate fixed point irrespective of architecture, modality, or ensemble structure.
Our work formalises these observations into a rigorous proof of convergence under the vanishing-exogenous-signal condition ($\alpha_{t}\to 0$), generalising the argument to any generative model trained via distributional objectives. Empirical studies indicate that collapse depends critically on the proportion of real versus synthetic data, with degradation emerging when synthetic data dominate and stability maintained when sufficient external signal persists (2). Our analysis provides the formal condition under which such degradation becomes inevitable.
4 Self-Referential Training as a Dynamical System
To analyse the dynamics of recursive self-improvement, we must first establish a formal mathematical framework. We generalise beyond specific architectures to any generative model that learns a probability distribution over a data space.
Definition 1 (Generative Model and Data Distributions).
Let $\mathcal{X}$ be a measurable data space.
  • Let $P$ be the true data distribution over $\mathcal{X}$, representing authentic, high-fidelity data (e.g., all human-generated text and images).
  • A generative model at iteration $t$, denoted $M_{t}$, is characterised by a probability distribution $Q_{t}$ over $\mathcal{X}$, which it can sample from. We assume $Q_{t}$ belongs to some family of distributions $\mathcal{Q}$ representable by the model architecture.
  • We assume an initial model $M_{0}$ with distribution $Q_{0}$ has been trained on samples drawn exclusively from $P$, such that $Q_{0}\approx P$. Due to finite data and model capacity, $Q_{0}\neq P$.
Definition 2 (Self-Referential Training Loop).
The transition from model $M_{t}$ to $M_{t+1}$ occurs via training on a mixture of exogenous and endogenous signal.
  • Let $\alpha_{t}\in[0,1]$ be the proportion of exogenous training signal at iteration $t$, meaning signal whose correctness or generating process is determined independently of the model’s current distribution. This may include fresh samples drawn from $P$, environmental feedback, simulator outcomes, or outputs validated by an external verifier.
  • The remaining proportion, $1-\alpha_{t}$, consists of endogenous samples generated from the current model distribution $Q_{t}$ (or an ensemble derived from it).
  • The training distribution for the next model, $P_{t}^{\prime}$, is the convex combination:
    $P_{t}^{\prime}=\alpha_{t}P+(1-\alpha_{t})Q_{t}.$ (9)
  • The new model $M_{t+1}$ is obtained by optimising its parameters to minimise divergence between its distribution $Q_{t+1}$ and the training distribution $P_{t}^{\prime}$. A common objective is to minimise KL divergence:
    $Q_{t+1}=\arg\min_{Q\in\mathcal{Q}}D_{KL}(P_{t}^{\prime}\,||\,Q).$ (10)
    This is equivalent to maximising the log-likelihood of data sampled from $P_{t}^{\prime}$.
The Singularity hypothesis of recursive self-improvement corresponds to the case where this iterative process, predominantly driven by self-generated data (i.e., $\alpha_{t}\to 0$), leads to a sequence of distributions $\{Q_{t}\}$ that become progressively “better” or “more intelligent”. Our thesis is that this process instead leads to convergence towards a degenerate fixed point $Q^{*}$, where $D_{KL}(P\,||\,Q^{*})>0$ and, critically, the entropy and diversity of $Q^{*}$ are lower than that of $P$.
4.1 Generalisation to Different Architectures
This framework is deliberately abstract to encompass various GenAI paradigms:
  • For LLMs, $\mathcal{X}$ is the space of token sequences and $Q_{t}$ is an autoregressive model. Self-referential training means fine-tuning the LLM on a mix of human text ($P$) and text generated by the LLM itself ($Q_{t}$).
  • For Diffusion Models, $\mathcal{X}$ is the space of images. The model learns a denoising function. Training on synthetic data means using images generated by the model in a previous iteration as clean examples for a new training run. The update rule (10) corresponds to re-optimising the denoising network on samples from the mixture distribution $P_{t}^{\prime}$.
  • For adversarial setups, such as GANs or models where a classifier provides a reward signal, the core logic holds. If the generator (the core generative component) is rewarded based on a classifier’s judgement, and that classifier is itself trained or fine-tuned on the generator’s outputs, a closed loop is formed. The ensemble’s knowledge becomes untethered from the external reality defined by $P$, leading to a similar collapse.
4.2 Proof of Convergence to a Degenerate State
The update rule in Equation (10) defines a discrete-time dynamical system on the space of probability distributions. Let $T_{\alpha}$ be the operator that maps $Q_{t}$ to $Q_{t+1}$ for a fixed $\alpha_{t}=\alpha$:
$Q_{t+1}=T_{\alpha}(Q_{t})=\arg\min_{Q\in\mathcal{Q}}D_{KL}(\alpha P+(1-\alpha)Q_{t}\,||\,Q)$ (11)
If the model family $\mathcal{Q}$ is sufficiently expressive to represent any mixture (e.g., if $\mathcal{Q}$ is the set of all probability distributions), then the solution is simply $Q_{t+1}=\alpha P+(1-\alpha)Q_{t}$. This is an exponentially weighted moving average, and its dynamics are straightforward.
Proposition 1 (Convergence of the Idealised Update Rule).
Assuming the model family $\mathcal{Q}$ has infinite capacity (i.e., can represent any distribution in the simplex of $\mathcal{X}$), if $Q_{t+1}=\alpha P+(1-\alpha)Q_{t}$ with a constant $\alpha\in(0,1]$, the sequence of distributions $\{Q_{t}\}_{t=0}^{\infty}$ converges to the true distribution $P$.
Proof.
By recursively expanding the update rule, we have:
$Q_{t}$ $=\alpha P+(1-\alpha)Q_{t-1}$
$=\alpha P+(1-\alpha)(\alpha P+(1-\alpha)Q_{t-2})$
$=\alpha P\sum_{k=0}^{t-1}(1-\alpha)^{k}+(1-\alpha)^{t}Q_{0}$
$=\alpha P\frac{1-(1-\alpha)^{t}}{1-(1-\alpha)}+(1-\alpha)^{t}Q_{0}$
$=(1-(1-\alpha)^{t})P+(1-\alpha)^{t}Q_{0}$
As $t\to\infty$, since $\alpha\in(0,1]$, we have $(1-\alpha)^{t}\to 0$. Therefore, $\lim_{t\to\infty}Q_{t}=P$. Remark: This result relies entirely on the assumption that $Q_{t+1}$ can perfectly capture the mixture. In reality, finite capacity introduces an approximation error $\delta$ at every step, which accumulates when $\alpha$ is small, as shown in Theorem 2. ∎
4.3 The Case of Pure Self-Reference ($\alpha=0$)
Let us now consider the crucial case for the Singularity hypothesis: a system that improves by learning exclusively from its own output. This corresponds to setting $\alpha=0$.
Theorem 2 (Entropy Decay in Closed-Loop Training).
Let the training dataset $\mathcal{D}_{t}$ at iteration $t$ be a finite set of $N$ samples drawn from $Q_{t}$. Let $Q_{t+1}$ be the empirical risk minimiser over $\mathcal{D}_{t}$. In the absence of external ground truth ($\alpha=0$) and assuming the model family $\mathcal{Q}$ has sufficient capacity to overfit, the differential entropy of the model sequence decreases in expectation:
$\mathbb{E}[H(Q_{t+1})]\leq H(Q_{t})-\Delta(N),$ (12)
where $\Delta(N)>0$ is a strictly positive term representing information loss due to finite sampling and the discrete approximation of continuous or high-dimensional supports. Consequently, $Q_{t}$ converges to a minimal-entropy distribution (a point mass or subset of modes) as $t\to\infty$.
Proof.
Let $\widehat{Q}_{t}$ be the empirical distribution formed by $N$ samples drawn from $Q_{t}$. By the properties of sampling from high-dimensional distributions, the support of $\widehat{Q}_{t}$ is a sparse subset of the support of $Q_{t}$. Specifically, for any distribution with tails (non-compact support), the probability that the finite sample support covers the true support is zero.
The model update $Q_{t+1}$ minimises $D_{KL}(\widehat{Q}_{t}\|Q)$. This is equivalent to maximising likelihood on the finite sample set. While standard maximum likelihood estimation is asymptotically unbiased, it exhibits finite-sample variance that manifests as overfitting to the sampled modes. The “missed” modes in $\widehat{Q}_{t}$ (events with probability $p<1/N$) are effectively assigned zero probability mass in the empirical target.
Critically, $H(\widehat{Q}_{t})<H(Q_{t})$ for finite $N$ due to the discretisation of the sample space. Since $Q_{t+1}$ is optimized to approximate $\widehat{Q}_{t}$, the sequence of entropies forms a supermartingale: $\mathbb{E}[H(Q_{t+1})\mid Q_{t}]\leq H(Q_{t})$. By the Martingale Convergence Theorem, $H(Q_{t})$ converges almost surely to a random variable with minimal entropy consistent with the fixed points of the update operator (i.e., mode collapse). ∎
This proves that in a closed loop, no growth in ‘intelligence’ is possible as no new knowledge is generated. The system is information-theoretically closed. This can also be seen from the perspective of the Data Processing Inequality (7).
Corollary 3 (Information-Theoretic Stagnation).
The self-referential training loop cannot increase the mutual information with the true distribution $P$.
Proof.
Consider the Markov chain $P\to Q_{t}\to Q_{t+1}$. The data processing inequality states that for any Markov chain $X\to Y\to Z$, we have $I(X;Z)\leq I(X;Y)$. In our case, this means $I(P;Q_{t+1})\leq I(P;Q_{t})$. The mutual information between the model’s state and the true state of the world can only decrease or stay the same with each iteration of self-training. Any imperfection in $Q_{t}$ (i.e., information about $P$ that $Q_{t}$ has lost) cannot be recovered by training on samples from $Q_{t}$. ∎
4.4 The Realistic Case ($\alpha_{t}\to 0$)
The most realistic scenario for a purported Singularity is one where an AI starts with access to human data but gradually becomes more autonomous, causing the proportion of authentic data $\alpha_{t}$ to approach zero over time. We will now show that this leads to model collapse.
Theorem 4 (Variance Amplification and Mean Shift).
Let the true distribution be $P$ and the model update be $Q_{t+1}=\mathcal{T}(P^{\prime}_{t})+\epsilon_{t}$, where $\mathcal{T}$ is the ideal update operator and $\epsilon_{t}$ is an approximation error term (due to SGD noise and finite sampling) with variance $\sigma^{2}_{\epsilon}$. If $\alpha_{t}\to 0$, the squared error between the model mean $\mu_{t}=\mathbb{E}_{x\sim Q_{t}}[x]$ and the true mean $\mu_{P}$ diverges or follows a random walk bounded only by the support diameter.
Proof.
Consider the mean of the distribution $\mu_{t}$. The update rule for the mixture $P^{\prime}_{t}=\alpha_{t}P+(1-\alpha_{t})Q_{t}$ implies the target mean is $\mu^{\prime}_{t}=\alpha_{t}\mu_{P}+(1-\alpha_{t})\mu_{t}$. The new model learns this mean with some error: $\mu_{t+1}=\mu^{\prime}_{t}+\xi_{t}$, where $\xi_{t}$ is a noise term corresponding to $\epsilon_{t}$. Substituting the target mean:
$\mu_{t+1}=(1-\alpha_{t})\mu_{t}+\alpha_{t}\mu_{P}+\xi_{t}.$ (13)
This describes an autoregressive process AR(1). As $\alpha_{t}\to 0$, the autoregressive coefficient $(1-\alpha_{t})\to 1$. The process approaches a random walk: $\mu_{t+1}\approx\mu_{t}+\xi_{t}$. Unlike the case where $\alpha>0$ (which provides a restoring force pulling $\mu_{t}$ back to $\mu_{P}$), the condition $\alpha_{t}\to 0$ removes the restoring force. The variance of the mean $\text{Var}(\mu_{t})$ grows linearly with $t$ in the random walk regime. Thus, the model distribution centre drifts away from the true distribution centre purely due to accumulated stochastic errors, confirming model collapse not just as mode-dropping, but as distributional drift. ∎
Theorem 5 (Scope Condition: Collapse Requires Vanishing Exogenous Signal).
Let $\{Q_{t}\}$ evolve under the update rule
$Q_{t+1}=\arg\min_{Q\in\mathcal{Q}}D_{KL}(P_{t}^{\prime}\,||\,Q),\quad P_{t}^{\prime}=\alpha_{t}P+(1-\alpha_{t})Q_{t}.$
If $\inf_{t}\alpha_{t}>0$ (i.e. the exogenous signal remains persistently bounded away from zero), and the model family $\mathcal{Q}$ has sufficient capacity, then in the idealised infinite-capacity limit the sequence $\{Q_{t}\}$ converges to $P$.
Conversely, the degenerative dynamics proven in Theorems 2 and the variance amplification result require $\alpha_{t}\to 0$.
Proof.
If there exists $\alpha_{*}>0$ such that $\alpha_{t}\geq\alpha_{*}$ for all $t$, then the update operator contains a persistent contraction term toward $P$. Under the infinite-capacity idealisation, the recursion reduces to
$Q_{t+1}=\alpha_{t}P+(1-\alpha_{t})Q_{t},$
which forms a weighted average with a non-vanishing restoring force. Standard contraction arguments imply convergence to $P$. The collapse results require that this restoring force vanish asymptotically, i.e. $\alpha_{t}\to 0$, in which case the system becomes asymptotically self-referential. ∎
The parameter $\alpha_{t}\in[0,1]$ denotes the proportion of exogenous training signal available at iteration $t$, that is, signal whose truth value or generating process is determined independently of the model’s current distribution. In the simplest instantiation used in this paper, this exogenous component is represented by fresh samples from the true distribution $P$.
The condition $\alpha_{t}\to 0$ means $\lim_{t\to\infty}\alpha_{t}=0$, i.e. for every $\varepsilon>0$ there exists $T$ such that $\alpha_{t}<\varepsilon$ for all $t\geq T$. Equivalently, the training mixture
$P^{\prime}_{t}=\alpha_{t}P+(1-\alpha_{t})Q_{t}$
becomes asymptotically self-referential, with the influence of exogenous signal vanishing in the limit.
This regime formalises the autonomy requirement implicit in certain Singularity-style recursive self-improvement arguments: the system must eventually rely predominantly on its own outputs rather than on a persistent external evaluator, simulator, or data-generating process. The scope clarifications in Section 4.9 make explicit that externally anchored self-improvement (fixed axioms and invariant evaluators) does not satisfy this vanishing-exogenous-signal assumption, even when data are produced via self-play or simulation.
Remark 6 (The Limits of Reinforcement Learning and Verifiers).
It is often argued that Reinforcement Learning (RL) with a verifier (e.g., a game engine or compiler) allows for self-improvement without external data ($\alpha=0$), as seen in systems like AlphaZero. However, this relies on the verifier providing a perfect, infinite-precision ground truth signal. In the context of AGI and LLMs operating in open-ended domains (language, reasoning, physical reality), no such perfect verifier exists.
If the verifier is itself a learned model (e.g., a Reward Model in RLHF), it is subject to the same collapse dynamics described in Theorem 2. If the verifier is a static proxy (e.g., string matching or simple heuristics), the model will exploit the metric (Goodhart’s Law), leading to semantic collapse, where the model optimises the proxy to the detriment of the underlying complexity. Thus, the impossibility results hold for any system where the ground truth is not strictly encoded in a formal, executable environment.
Lemma 7 (Asymptotic self-reference of the training distribution).
If $\alpha_{t}\to 0$, then for any fixed $t$ the mixture distribution satisfies
$\|P^{\prime}_{t}-Q_{t}\|_{\mathrm{TV}}\;\leq\;\alpha_{t}\,\|P-Q_{t}\|_{\mathrm{TV}}\;\leq\;\alpha_{t},$
and hence $\|P^{\prime}_{t}-Q_{t}\|_{\mathrm{TV}}\to 0$ as $t\to\infty$, where $\|\cdot\|_{\mathrm{TV}}$ denotes total variation distance.
Proof.
By linearity of mixtures, $P^{\prime}_{t}-Q_{t}=\alpha_{t}(P-Q_{t})$, so $\|P^{\prime}_{t}-Q_{t}\|_{\mathrm{TV}}=\alpha_{t}\|P-Q_{t}\|_{\mathrm{TV}}\leq\alpha_{t}$ since $\|P-Q_{t}\|_{\mathrm{TV}}\leq 1$ for probability measures. Taking $t\to\infty$ and using $\alpha_{t}\to 0$ completes the proof. ∎
4.5 Extension to Multi-Modal Ensembles
One might think that an ensemble of multi-modal models could prevent collapse. This is true until the new training data set is exhausted. Let us formalise this scenario and show that it eventually suffers the same fate.
Definition 3 (Multi-Modal Training).
Consider an ensemble of $N$ models including possible different data types $\{M_{1},\dots,M_{N}\}$ with corresponding distributions $\{Q_{t}^{1},\dots,Q_{t}^{N}\}$. The synthetic data are drawn from a mixture of these models:
$R_{t}=\sum_{i=1}^{N}\omega_{i}Q_{t}^{i},\quad\text{where }\sum\omega_{i}=1,\omega_{i}\geq 0.$ (14)
Each model $M_{j}$ is then trained on a distribution $P_{t,j}^{\prime}=\alpha_{t}P+(1-\alpha_{t})R_{t}$.
Theorem 8 (Collapse of Multi-Model Ensembles).
As $\alpha_{t}\to 0$, the ensemble of models converges to a consensus fixed-point distribution $R^{*}\neq P$. The individual distributions $\{Q_{i}^{*}\}$ may not be identical, but the mixture they form will be a stationary, degenerate distribution.
Proof.
As $\alpha_{t}\to 0$, each model $M_{j}$ is being trained to imitate the ensemble’s average distribution from the previous step, $R_{t-1}$.
$Q_{t}^{j}\approx\arg\min_{Q\in\mathcal{Q}_{j}}D_{KL}(R_{t-1}\,||\,Q)$ (15)
The entire system’s state is now described by the mixture distribution $R_{t}$. The update rule for the mixture is $R_{t}\approx\sum\omega_{i}\arg\min_{Q\in\mathcal{Q}_{i}}D_{KL}(R_{t-1}\,||\,Q)$. This is still a closed loop. The ensemble is learning from itself. Any shared biases or errors across the models will be reinforced. Any aspect of the true distribution $P$ that is collectively under-represented by the initial ensemble $R_{0}$ will be progressively forgotten as $\alpha_{t}\to 0$. The system converges to a fixed-point mixture $R^{*}$ that represents the consensus reality of the initial models, not the true reality $P$. Diversity may be lost more slowly than in a single-model case, but the fundamental information-theoretic barrier remains: no new information about $P$ can be created from within the closed system. ∎
4.6 Closed-Loop Training vs Externally Anchored Optimisation
The model collapse results presented above apply to a specific structural regime. It is therefore essential to distinguish between two fundamentally different forms of recursive training.
1. Closed-Loop Density Matching.
In the collapse setting, the model is iteratively trained to approximate a distribution that increasingly consists of its own samples. When $\alpha_{t}\to 0$, the update becomes:
$Q_{t+1}\approx\arg\min_{Q}\mathbb{E}_{x\sim Q_{t}}[-\log Q(x)].$
This defines a self-referential transformation on the space of distributions.
In this regime:
  • The objective is implicit in the data distribution itself.
  • There is no external correctness criterion beyond likelihood.
  • The optimisation landscape is shaped by $Q_{t}$ rather than by an independently specified target.
The collapse theorem characterises precisely this regime.
2. Externally Anchored Optimisation.
By contrast, many canonical examples of “self-improvement” operate under additional structural assumptions:
  1. 1.
    A fixed axiomatic base (e.g., game rules or physical laws).
  2. 2.
    An externally defined objective or verifier.
  3. 3.
    A bounded task domain in which improvement is measured.
In such systems, even if training data are generated through self-play or simulation, the evaluation signal is not defined by the model’s own distribution.
Instead, the update operator takes the form:
$Q_{t+1}=\mathcal{T}(Q_{t};\mathcal{E}),$
where $\mathcal{E}$ denotes an external evaluator or environment.
The asymptotic behaviour of this operator is governed by properties of the external objective, not by self-consistency alone.
The collapse results apply to closed-loop generative retraining without persistent exogenous signal. They do not automatically extend to systems in which optimisation remains externally grounded.
This distinction is crucial when drawing conclusions about AGI or recursive self-improvement.
5 Implications for AGI, ASI and Singularity Arguments
The results established in this paper apply to a specific structural regime: recursive generative retraining in which the proportion of exogenous signal $\alpha_{t}$ vanishes asymptotically.
Under this regime, the system becomes information-theoretically closed. Learning reduces to self-consistency under a distributional objective, and the update operator converges to a fixed point determined by its own prior state. Theorems 2 and the variance amplification result show that in this regime entropy decays and distributional drift accumulates.
The theorems establish that:
  • Purely self-referential density matching ($\alpha_{t}\to 0$) induces contraction toward a degenerate or distorted stationary distribution.
  • Mutual information with the true generating distribution $P$ cannot increase under distribution-only updates.
  • Autonomy in the strong sense of vanishing external grounding leads to asymptotic self-reference rather than unbounded improvement.
The results do not prove that all forms of recursive self-improvement collapse.
If $\inf_{t}\alpha_{t}>0$, meaning the system receives persistent exogenous signal, then the contraction toward $P$ remains active. Systems operating under fixed axioms, externally defined objectives, or invariant verifiers (e.g. formally specified environments) do not satisfy the $\alpha_{t}\to 0$ condition.
Such systems may converge to task-optimal stationary distributions rather than degenerate measures. This distinction is consistent with classical results in stochastic approximation theory, where convergence under persistent excitation is guaranteed when the update receives non-vanishing external signal (1). The collapse theorems proven here apply specifically to the complementary regime in which the excitation term vanishes asymptotically ($\alpha_{t}\to 0$).
Therefore, any claim that a distributional generative model will achieve unbounded recursive improvement must specify whether its learning signal becomes endogenised. If the system asymptotically relies on its own outputs without persistent grounding, collapse follows under the assumptions of this paper. If improvement instead remains externally anchored, the asymptotic regime is fundamentally different and must be analysed under persistent-excitation dynamics.
The Singularity hypothesis, when framed as fully autonomous recursive density matching, is mathematically incompatible with sustained growth under current KL-based generative learning paradigms.
5.1 Causality and Neurosymbolic Approaches as a Path Beyond Collapse
The results presented show that recursive self-training with objectives grounded in Kullback–Leibler (KL) divergence leads to model collapse. It is important to emphasise that KL divergence is not an arbitrary choice: it underlies nearly all loss functions used in current deep learning, particularly in the transformer family of models. Cross-entropy minimisation is equivalent to minimising $D_{KL}(P\,||\,Q)$, and maximum likelihood estimation also reduces to this divergence under common assumptions. Thus, our analysis captures the essential training dynamics of statistical deep learning without loss of generality.
Nearly all practical loss functions used in Large Language Models (LLMs) and related deep learning architectures reduce to KL divergence or very close relatives. This establishes that our collapse results apply to the entire class of current statistical deep learning approaches without loss of generality.
  • Cross-Entropy Loss. The token-level cross-entropy used in transformers is exactly equivalent to the KL divergence between the true data distribution $P$ and the model distribution $Q$. Specifically,
    $H(P,Q)=H(P)+D_{KL}(P\,||\,Q),$
    where $H(P)$ is constant with respect to $Q$. Minimising cross-entropy is therefore equivalent to minimising KL divergence.
  • Mean Squared Error (MSE). In regression, MSE minimisation is equivalent to maximum likelihood estimation under Gaussian noise assumptions. Maximum likelihood corresponds to minimising the KL divergence between the empirical data distribution and the Gaussian model distribution, making MSE a special case of KL.
  • Binary Cross-Entropy / Logistic Loss. Logistic regression minimises the cross-entropy between Bernoulli distributions, which again reduces to KL divergence between $P$ (empirical labels) and $Q$ (model predictions).
  • Categorical Cross-Entropy. Multiclass classification with softmax uses the categorical cross-entropy, which is the KL divergence between a one-hot true label distribution and the predicted probability distribution.
  • Other Variants. Many widely used losses (e.g., focal loss, label smoothing) are weighted or regularised versions of cross-entropy, and thus still within the KL framework. Generative Adversarial Networks (GANs) minimise Jensen–Shannon divergence, which belongs to the broader $f$-divergence family of which KL is a canonical member. Even Wasserstein distances, while distinct, are less commonly used in LLM training and do not alter the essential correlation-based character of the optimisation.
Thus, KL divergence provides a unifying mathematical framework that encompasses the most common objectives in deep learning. Our collapse proofs, therefore, apply broadly to existing transformer and generative model training regimes.
5.2 Statistical Loss Functions Contribute to Collapse
The reliance on KL means that optimisation is driven by correlation, not causation. Models trained under KL minimise distributional divergence but cannot infer or extrapolate underlying mechanisms. This explains why recursive training contracts distributions, amplifies biases, and converges to degenerate fixed points: KL training lacks a mechanism to preserve diversity beyond what is already present in the data mixture.
An alternative is to replace correlation-based optimisation with causal or algorithmic objectives. Instead of training future models as
$Q_{t+1}=\arg\min_{Q}D_{KL}(P_{t}^{\prime}\,||\,Q),$
we could envisage
$Q_{t+1}=\arg\min_{Q}D_{\text{causal}}(M(P_{t}^{\prime}),M(Q)),$
where $M(\cdot)$ maps raw distributions to causal models, e.g., structural causal graphs. Optimising causal divergence aligns models with invariances under intervention, not just statistical regularities. This allows even synthetic data to yield new knowledge if it encodes counterfactual or interventional information.
Such neurosymbolic architectures, combining statistical learners with causal inference and algorithmic probability (as in CTM and BDM), could in principle break the collapse dynamics. Unlike KL-based systems, they are capable of producing genuinely synthetic knowledge, rather than endlessly recombining correlations. However, current LLMs remain confined to KL-like objectives and thus to the collapse trajectory we have proven.
If the incorporation of causal and algorithmic principles for the update rule is reformulated not as
$Q_{t+1}=\arg\min_{Q}D_{KL}(P_{t}^{\prime}\,||\,Q),$
but rather as
$Q_{t+1}=\arg\min_{Q}D_{\text{causal}}(M(P_{t}^{\prime}),M(Q)),$
where $M(\cdot)$ denotes a causal representation of a distribution, for example, a structural causal model or a symbolic regression (26; 28), then the system optimises not for correlation but for invariance under intervention. In this framework, even synthetic data can yield new information if they encode counterfactual or interventional predictions. Unlike KL-based training, which contracts toward fixed points, causal objectives permit the creation of genuinely synthetic knowledge, in line with Kant’s distinction between analytic and synthetic judgments.
5.3 Causal correction and rate of (non-)collapse
We now formalise how the rate of convergence to collapse (or escape from it) depends on the quality of (i) the causal component and (ii) the symbolic component in a neurosymbolic pipeline. Throughout, $P$ denotes the target (real-world) distribution, $Q_{t}$ the model distribution after iteration $t$, and $\alpha_{t}\in[0,1]$ the fraction of fresh ($P$-drawn) data used at iteration $t$. We work with an $f$-divergence $D_{f}(\cdot\|\cdot)$ that includes $\mathrm{KL}$ as a special case; all arguments below therefore cover the standard cross-entropy/negative log-likelihood training used by LLMs and most deep models.
Operators.
Let the one-step update be a composition of three operators acting on distributions:
$Q_{t+1}\;=\;\underbrace{\mathcal{T}_{\alpha_{t}}}_{\text{statistical learner}}\;\circ\;\underbrace{\mathcal{C}_{t}}_{\text{causal correction}}\;\circ\;\underbrace{\Pi_{\mathcal{S}}}_{\text{symbolic projection}}\,(Q_{t}),$
where:
  1. 1.
    $\mathcal{T}_{\alpha_{t}}$ is the statistical (likelihood) update to fit the mixture $P^{\prime}_{t}=\alpha_{t}P+(1-\alpha_{t})Q_{t}$ (Sec. 3). In the well-specified idealisation one has $\mathcal{T}_{\alpha_{t}}(Q)=\arg\min_{Q^{\prime}}D_{\mathrm{KL}}(P^{\prime}_{t}\|Q^{\prime})$.
  2. 2.
    $\Pi_{\mathcal{S}}$ is a projection onto a feasible set $\mathcal{S}$ of symbolic/axiomatic constraints (e.g., conservation, monotonicity, type rules, physical/biological invariants). Formally, $\Pi_{\mathcal{S}}(Q)=\arg\min_{R\in\mathcal{S}}D_{f}(R\|Q)$.
  3. 3.
    $\mathcal{C}_{t}$ is a causal-correction operator built from interventional queries; e.g., it replaces or reweights some conditionals $Q(X\mid\mathrm{do}(Z))$ using estimates derived from a causal model $\mathcal{G}$ and interventional data. We quantify its strength by how much it contracts the divergence to $P$ on the parts it corrects.
Quantifying symbolic and causal “power”.
We define two per-iteration contraction factors:
(Symbolic) $D_{f}\!\left(P\;\middle\|\;\Pi_{\mathcal{S}}(Q)\right)\;\leq\;\underbrace{\sigma}_{\in(0,1]}\;D_{f}\!\left(P\;\middle\|\;Q\right)\;+\;\delta_{s},$ (16)
(Causal) $D_{f}\!\left(P\;\middle\|\;\mathcal{C}_{t}(R)\right)\;\leq\;\underbrace{\kappa_{t}}_{\in(0,1]}\;D_{f}\!\left(P\;\middle\|\;R\right)\;+\;\delta_{c,t},$ (17)
where $0<\sigma\leq 1$ captures the strength of symbolic constraints (smaller is better), and $0<\kappa_{t}\leq 1$ captures the effective causal correction at iteration $t$ (again, smaller is better). The additive terms $\delta_{s},\delta_{c,t}\geq 0$ capture model/estimation imperfections (finite data, solver tolerance, partial coverage of the graph, etc.). In practice,
$\kappa_{t}\;=\;1-\eta_{c}\,\phi_{t},\qquad\phi_{t}\in[0,1]$
where $\eta_{c}\in(0,1]$ encodes identification strength of the causal queries (e.g., instrument strength, overlap, do-operator availability), and $\phi_{t}$ is the coverage of interventional updates at step $t$ (fraction of conditionals actually corrected).
The contraction factor $\sigma<1$ is not arbitrary but is a direct consequence of the Coding Theorem. Let $\mathcal{H}$ be the total hypothesis space of size $|\mathcal{H}|$. A purely statistical learner effectively searches this entire space for correlations. A symbolic learner projects onto a subspace $\mathcal{S}\subset\mathcal{H}$ defined by programs $p$ with Kolmogorov complexity $K(p)\leq L$. By the Coding Theorem, the probability mass is concentrated on simple programs: $m(x)\approx 2^{-K(x)}$. If the true mechanism $M$ has low complexity ($M\in\mathcal{S}$), the projection $\Pi_{\mathcal{S}}$ eliminates the vast majority of the search space (high-complexity noise and overfitting candidates). The reduction in the volume of the search space corresponds to an information gain of approximately $|\mathcal{H}|-|\mathcal{S}|$ bits. In terms of divergence, this forces the model distribution $Q$ to align with the algorithmic structure of $P$, ensuring that
$D_{f}(P\|\Pi_{\mathcal{S}}(Q))\ll D_{f}(P\|Q),$
which implies an effective contraction $\sigma\ll 1$ whenever the data is generated by a computable process.
The quantities $\sigma$ and $\kappa_{t}$ in Eqs. (16) and (17) are contraction factors. Their magnitude determines the rate at which iterative updates reduce discrepancy with respect to the target distribution or mechanism. Throughout, smaller values correspond to stronger correction.
Formally, consider a generic inequality of the form
$D_{f}^{(t+1)}\;\leq\;c\,D_{f}^{(t)}+\delta,\qquad c\in(0,1].$ (18)
Iterating (18) yields
$D_{f}^{(n)}\;\leq\;c^{n}D_{f}^{(0)}\;+\;\sum_{i=0}^{n-1}c^{i}\delta,$ (19)
which shows that the discrepancy decays exponentially at rate $c$ up to an error floor determined by $\delta$. When $c\ll 1$, convergence is fast; when $c\approx 1$, improvement per iteration is negligible; and when $c=1$, no contraction occurs.
In Eq. (16), the factor $\sigma$ quantifies the strength of symbolic constraints. The projection $\Pi_{\mathcal{S}}$ enforces syntactic, grammatical, or invariant structure, thereby eliminating large regions of the hypothesis space. A small value of $\sigma$ indicates that symbolic structure is highly informative, leading to a rapid collapse toward low-complexity representations. From an algorithmic information perspective, this corresponds to strong compression of admissible descriptions, consistent with CTM-based estimation.
In Eq. (17), the factor $\kappa_{t}$ measures the effective causal correction at iteration $t$. Writing
$\kappa_{t}=1-\eta_{c}\phi_{t},$
makes explicit that causal contraction improves with both the identification strength $\eta_{c}$ of interventions or perturbations and their coverage $\phi_{t}$. Smaller values of $\kappa_{t}$ indicate that causal updates eliminate large classes of incompatible mechanisms per iteration, reflecting strong perturbation coherence and high causal informativeness.
By contrast, the statistical update in Eq. (20) induces contraction only in distribution space. Although the factor $(1-\alpha_{t})$ also yields contraction in the sense of Eq. (19), this improvement is limited by the data-processing inequality and cannot increase information about the underlying generative mechanism. Statistical contraction therefore represents a weak, degenerate regime in which learning reduces to smoothing or reweighting existing distributional information.
Taken together, these observations establish a hierarchy. Statistical updates contract discrepancies between distributions but cannot recover mechanisms. Symbolic updates contract hypothesis space more aggressively by exploiting structural constraints. Causal updates operating in program space provide the strongest contraction by directly eliminating incompatible generative mechanisms. Smaller contraction factors thus correspond to stronger explanatory and corrective power per iteration.
Statistical learner under synthetic drift.
Unlike the symbolic and causal updates above, this step does not perform causal correction. It characterises the behaviour of a learner restricted to distributional information only, as arises under synthetic drift or when algorithmic structure cannot be accessed.
For the likelihood update on $P^{\prime}_{t}=\alpha_{t}P+(1-\alpha_{t})Q_{t}$, standard information-geometry arguments yield, for common choices of $D_{f}$ including $\mathrm{KL}$:
$D_{f}\!\left(P\;\middle\|\;\mathcal{T}_{\alpha_{t}}(S)\right)\;\leq\;\bigl(1-\alpha_{t}\bigr)\;D_{f}\!\left(P\;\middle\|\;S\right)\;+\;\delta_{\mathrm{stat},t},$ (20)
where $\delta_{\mathrm{stat},t}\to 0$ in the well-specified, infinite-data limit; otherwise it scales with capacity/optimisation error. In other words, even when statistical learning converges optimally, it may converge to a causally meaningless solution.
Proposition 9 (DPI bound for distribution-only learning and the resulting causal hierarchy).
Let $\mathcal{A}$ be any distribution-only learner/estimator that maps an observed sample (or an empirical distribution) to a model distribution,
$\mathcal{A}:\ \widehat{P}\ \mapsto\ Q_{\theta(\widehat{P})}$,
and suppose the learning pipeline factors through a Markov chain of the form
$M\;\longrightarrow\;X\;\longrightarrow\;\widehat{P}(X)\;\longrightarrow\;\theta(\widehat{P})\;\longrightarrow\;Q_{\theta},$
where $M$ denotes the (unknown) data-generating mechanism and $X$ the observed data. Then the mutual information between the mechanism and the learned model is bounded by the data-processing inequality:
$I(M;Q_{\theta})\;\leq\;I(M;X).$ (21)
In particular, no distribution-only learning update—including one driven by minimising $\mathrm{KL}$ or any $f$-divergence—can increase the information the learner has about the generating mechanism beyond what is already contained in the observations.
Moreover, if the learner’s update further compresses $X$ into a finite-dimensional summary (e.g. empirical moments, sufficient statistics under a misspecified family, or a finite-capacity parametric model), then typically
$I(M;Q_{\theta})\;<\;I(M;X),$ (22)
with strict inequality whenever the map $X\mapsto Q_{\theta}$ is non-invertible on the support induced by $M$.
Proof.
The assumed pipeline defines the Markov chain
$M\to X\to Q_{\theta},$
[…]
$I(M;Q_{\theta})\leq I(M;X),$
which proves (21). If $X\mapsto Q_{\theta}$ is not one-to-one (equivalently, if there exist distinct $x\neq x^{\prime}$ with $Q_{\theta(x)}=Q_{\theta(x^{\prime})}$ on a set of non-zero probability under the joint distribution induced by $M$), then information is lost and the inequality is strict, yielding (22). ∎
Proposition 9 makes the hierarchy in this paper explicit: purely statistical updates contract distances between distributions (e.g. via $\mathrm{KL}$) but cannot increase information about mechanisms. Symbolic and causal updates that operate in program/mechanism space are not constrained in the same way, because they do not merely post-process $X$; they introduce additional mechanistic hypotheses and perturbation-coherence constraints that are not representable as a distribution-only map.
5.4 From Statistical to Algorithmic Integration
We now formalise how CTM and BDM allow us to transcend the limitations of statistical learning derived in Section 5.5.
Proposition 10 (Escaping DPI via Universal Priors).
While purely statistical learning is bound by the Data Processing Inequality (DPI) such that $I(M;Q_{t+1})\leq I(M;Q_{t})$, a generative programme synthesis approach $\mathcal{G}$ can effectively increase the mutual information with the true mechanism $M$ by conditioning on the Universal Distribution $m$.
Proof.
The DPI constraint $I(X;Z)\leq I(X;Y)$ for a Markov chain $X\to Y\to Z$ assumes that $Z$ is computed solely from $Y$. In Algorithmic Information Dynamics, the update is $(Q_{t},\mathcal{U})\to Q_{t+1}$, where $\mathcal{U}$ is an enumeration of Turing Machines ordered by length $2^{-|p|}$. The operator $\mathcal{G}$ selects programmes $p\in\mathcal{U}$ that reproduce the data $X$ within error $\epsilon$. This introduces exogenous information (Occam’s bias). If the true mechanism $M$ is simple, the prior $m(p)$ assigns high probability to $M$. The statistical learner assigns mass based on frequency (correlation); the algorithmic learner assigns mass based on descriptional complexity (causation). When data $X$ is sparse (the collapse regime), the intersection of data constraints and the algorithmic prior $m(p)$ can uniquely identify $M$, recovering information hidden to statistical projections. ∎
We have established that purely statistical updates lead to entropy decay (Theorem 2) and variance drift (Theorem 3). We now detail how the tools defined above counteract these specific failure modes.
Theorem 2 proves that models trained on finite samples $\widehat{Q}_{t}$ lose diversity because the “tails” are statistically invisible. Algorithmic Probability solves this by generative implication. Let the observed data be $x$. An algorithmic learner searches for the minimal program $p^{*}$ such that $U(p^{*})=x$. If $x$ is generated by a lawful mechanism, $p^{*}$ implicitly defines the entire distribution, including the tails missing from the sample. Formally, the algorithmic update expands support via the program’s domain:
$\text{supp}(Q_{alg})\;=\;\{y\mid\exists p,U(p)=y,K(p)\approx K(p^{*})\}.$ (23)
Thus, CTM restores entropy by re-deriving the underlying law necessitating the unseen data.
5.4.1 Countering Variance Drift: The Symbolic Anchor
Theorem 3 showed that without external grounding ($\alpha\to 0$), the model mean follows a random walk. Symbolic constraints act as a discretisation anchor. While a continuous parameter vector can drift infinitesimally, a program cannot; it must change to a distinct valid program. The update rule becomes:
$p_{t+1}\;=\;\arg\min_{p}\left[-\log P(D|p)+\lambda K(p)\right].$ (24)
The term $K(p)$ acts as a potential barrier. Small statistical noise is insufficient to jump to the next simplest program, locking the model to the simplest explanation $p^{*}$.
The algorithmic update ensures that $I(M;Q_{t+1}^{alg})\geq I(M;Q_{t}^{stat})$, explicitly escaping the data-processing inequality (as per Proposition 10) by injecting the universal prior $m$. This allows the system to maintain alignment with the ground truth mechanism $M$ even when $\alpha_{t}\to 0$.
5.5 The Algorithmic Solution to Entropy Decay and Drift
We have established that purely statistical updates lead to entropy decay (Theorem 2) and variance drift (Theorem 3) because they rely exclusively on the empirical properties of finite samples. To halt and reverse these dynamics, we must move from statistical inference to algorithmic inference. We now detail exactly how the tools defined in the previous section (CTM and BDM) mechanistically counteract the specific failure modes derived in Section 4.
5.5.1 Countering Entropy Decay: Generative Implication
Theorem 2 proves that models trained on finite samples $\widehat{Q}_{t}$ lose diversity because the “tails” (rare events) are statistically invisible. Standard regularisation (e.g., weight decay) smooths the distribution but does not recover the specific structure of the lost tails.
Algorithmic Probability, approximated by CTM (Eq. 6), solves this by generative implication. Let the observed data be a sequence $x$. A statistical learner sees $x$ as a collection of frequencies. An algorithmic learner searches for the minimal program $p^{*}$ such that $U(p^{*})=x$. Crucially, if the data $x$ is generated by a lawful mechanism (e.g., the sequence $2,4,6,8$), the minimal program $p^{*}$ (“print even numbers”) implicitly defines the entire distribution, including the tails that were missing from the sample (e.g., $10,12,\dots$).
Formally, while the statistical update shrinks support to the observed samples, $\text{supp}(Q_{stat})\subseteq\text{supp}(\widehat{Q}_{t})$, the algorithmic update expands support via the program’s domain:
$\text{supp}(Q_{alg})\;=\;\{y\mid\exists p,U(p)=y,K(p)\approx K(p^{*})\}.$ (25)
Thus, CTM restores the lost entropy not by adding random noise, but by re-deriving the underlying law that necessitates the existence of the unseen data.
5.5.2 Countering Variance Drift: The Symbolic Anchor
Theorem 3 showed that without external grounding ($\alpha\to 0$), the model mean $\mu_{t}$ follows a random walk $\mu_{t+1}\approx\mu_{t}+\xi_{t}$ driven by optimisation noise.
Symbolic constraints (Section 5.4) and algorithmic complexity act as a discretisation anchor. The space of low-complexity programs is sparse. While a continuous parameter vector can drift infinitesimally ($\mu_{t}\to\mu_{t}+\epsilon$), a program cannot “drift” slightly; it must change to a distinct valid program. The update rule under algorithmic constraint becomes:
$p_{t+1}\;=\;\arg\min_{p}\left[-\log P(D|p)+\lambda K(p)\right].$ (26)
The term $K(p)$ acts as a significant potential barrier. Small statistical noise $\xi_{t}$ is insufficient to jump the gap to the next simplest program. Consequently, the model state is “locked” to the simplest explanation $p^{*}$ until significantly contradictory evidence accumulates. This quantisation of the hypothesis space prevents the continuous degradation (random walk) characteristic of purely neural updates.
5.5.3 Refining the Contraction
We can now refine the symbolic contraction inequality (Eq. 16). The factor $\sigma$ is not arbitrary but is derived from the density of computable objects. By the Coding Theorem, the probability mass of the hypothesis space concentrates on simple programs: $m(x)\approx 2^{-K(x)}$. When we project the learned distribution $Q_{t}$ onto the set of distributions generated by programs of length $L$, we effectively discard the high-complexity “noise” that constitutes the drift $\xi_{t}$.
The algorithmic update ensures that $I(M;Q_{t+1}^{alg})\geq I(M;Q_{t}^{stat})$, explicitly escaping the data-processing inequality (as per Proposition 10) by injecting the universal prior $m$. This allows the system to maintain alignment with the ground truth mechanism $M$ even when the external data fraction $\alpha_{t}\to 0$, provided that the underlying reality $M$ remains algorithmically simple.
6 Conclusion
We have formalised recursive self-training in generative models as a discrete-time dynamical system and derived two structural failure modes that arise when the proportion of exogenous signal $\alpha_{t}$ vanishes asymptotically. Under this autonomy condition, the update operator becomes increasingly self-referential. Entropy decays due to finite-sample contraction, and distributional drift accumulates through variance amplification. The system converges not toward expanding intelligence, but toward stationary fixed points determined by its own prior state.
The impossibility result is conditional rather than universal. It applies to recursive density matching under KL-type objectives when the influence of externally grounded signal disappears in the limit. If $\alpha_{t}\to 0$, the system becomes information-theoretically closed: no new information about the generating distribution can be created through self-sampling alone. Mutual information with the true distribution cannot increase, and optimisation reduces to self-consistency rather than improvement relative to the world.
This conditional structure is crucial for evaluating claims about AGI, ASI, and the Singularity. Strong versions of these hypotheses typically posit systems that become increasingly autonomous and require little to no human or environmental intervention in order to improve themselves. Such autonomy implicitly entails a regime in which exogenous signal becomes negligible relative to internally generated data. In that regime, our theorems apply directly. Recursive self-improvement framed as progressively self-contained generative retraining cannot yield unbounded growth under standard distributional learning dynamics.
By contrast, systems that retain persistent external grounding ($\inf_{t}\alpha_{t}>0$), operate under fixed axioms with invariant evaluators, or optimise against formally specified environments fall outside the collapse regime proven here. Their asymptotic behaviour must be analysed under persistent-excitation dynamics rather than closed-loop contraction. Our results therefore do not rule out improvement in externally anchored systems; they rule out fully autonomous recursive density matching as a path to indefinite intelligence growth.
The deeper limitation is structural. KL-based and related divergence objectives optimise correlations within observed distributions but do not recover generative mechanisms. When such objectives dominate the update rule, recursive retraining contracts distributional diversity and amplifies internal biases. To escape these attractor dynamics, learning must move beyond analytic correlation toward synthetic mechanism discovery.
Neurosymbolic integration grounded in algorithmic information theory provides a principled direction. By incorporating program synthesis, algorithmic probability, and causal perturbation analysis through tools such as CTM and BDM, the learning objective can shift from distributional fit to mechanistic coherence. In such a framework, improvement is measured not by lower perplexity alone but by convergence toward simpler, causally invariant generative programmes consistent with the data.
In summary, fully autonomous recursive generative retraining under standard statistical objectives leads to degenerative fixed points rather than intelligence explosion. Sustained self-improvement requires persistent grounding or a transition from distributional optimisation to mechanism-based inference. The Singularity, if it is to be coherent, cannot be built upon closed-loop density matching alone.